9.3A simple random sample of size n=24 is drawn from a population that is normally distributed. The sample mean is found to be x
Answer #1

Given that, n=24 w = 68 S-13 d. f =n-1 = 24-1 d. f =23 Confidence level 951 x = level of significance x=1-CL = 1-951.=1-0.95